OpenStudy (anonymous):

What is the inverse cotangent of (square root of 3) divided by (3)?

OpenStudy (anonymous):

\[\cot-1(\sqrt{3}/3) = \pi/3\]

OpenStudy (anonymous):

\[\cot^{-1}(\sqrt{3}/3) = \pi/3\] <-- Fixed your typo. In TeX the "^" symbol only takes the first number after it so you have to put the negative exponents in between braces like this: {-1}. You can right click on the equation to see the TeX formatting.